# FedMLAC: Mutual Learning Driven Heterogeneous Federated Audio Classification

### v1

@@ 5.3 Discussion
The convergence rate of FedMLAC is $\mathcal{O}(1/\sqrt{T})$ when $\eta\propto 1/\sqrt{T}$, which is standard for non-convex optimization in FL. The mutual learning mechanism aligns local and global objectives, reducing the impact of $\sigma^{2}$ by regularizing local updates with global knowledge. The LPA strategy ensures robustness against data poisoning by bounding $\delta^{2}$, as it prunes anomalous updates that could destabilize aggregation. The lightweight Plug-in model reduces communication overhead, enhancing scalability in heterogeneous environments.
@@ 6 Experiments
In this section, we conduct comprehensive experiments to evaluate the effectiveness of the proposed FedMLAC framework. Our evaluation spans four benchmark datasets covering both speech and non-speech audio classification tasks. We assess FedMLAC in terms of classification accuracy, robustness to noisy and heterogeneous data, and personalization capability, comparing it against three representative state-of-the-art FL baselines.
6.1 Experimental Settings
@@
Datasets. We evaluate FedMLAC on four publicly available datasets widely adopted in federated audio classification (FedAC), as summarized in Table 2. (1) Google Speech Commands (GSC) [68] is a large-scale benchmark for keyword spotting, comprising over 100,000 short utterances covering 35 command words. (2) Urban Sound [69] includes environmental audio samples spanning 10 everyday sound event classes, such as sirens, drilling, and dog barking. (3) IEMOCAP [70] is a multimodal dataset for emotion recognition. In this work, we focus on its audio-only component, selecting four commonly used emotions from the improvised sessions. (4) CREMA-D [71] features speech clips performed by actors under emotional prompts. We consider the audio channel and select four target emotions consistent with prior studies. To ensure fair comparison across datasets, we apply consistent preprocessing strategies including silence trimming, segmentation, and acoustic feature extraction (e.g., Mel spectrograms or pretrained APC embeddings), following the protocol proposed in [72].
Data Partition. To simulate realistic client distributions in federated settings, we adopt dataset-specific non-IID partitioning strategies. For the Google Speech Commands and CREMA-D datasets, we assign samples to clients based on speaker identity, ensuring that each client contains utterances from a single speaker. This setup reflects naturally occurring data silos in speech-based systems. In the case of IEMOCAP, we divide data by actor identity, with each client corresponding to the recordings of one actor. This preserves speaker consistency within clients while exposing variation across clients. For the Urban Sound dataset, we follow a label distribution skewing strategy using the Dirichlet allocation method, as commonly adopted in FL studies [73]. Specifically, we vary the Dirichlet concentration parameter $\alpha\in\{0.1,0.3,0.5,1.0\}$ to control the degree of data heterogeneity. Smaller values of $\alpha$ lead to greater data heterogeneity across clients, whereas larger values result in more balanced and uniform data distributions.
AC Models. To support AC in both homogeneous and heterogeneous FL settings, we construct a suite of Convolutional Recurrent Neural Network (CRNN) architectures with varying levels of complexity as listed in Table 3. All models share a consistent input-output format and are compatible with the same audio preprocessing pipeline (e.g., Mel-spectrogram or APC features), ensuring fair performance comparison across settings.
@@
Metric. We evaluate performance on the Google Speech Commands dataset using classification accuracy, averaged over five independent runs with different random seeds. For the other three datasets, we adopt the F1 score as the evaluation metric, reporting the mean and standard deviation across all runs. F1 score is used to account for class imbalance, which is common in emotion recognition and sound event classification tasks.
Baselines. We compare our method with three widely used FL algorithms: FedAvg [10], FedProx [61], and FedOPT [62]. These methods serve as strong general-purpose baselines and are commonly adopted across diverse FL tasks, including audio applications. We do not include FedAC-specific approaches due to the lack of standardized implementations and their reliance on incompatible datasets or architectures, which hinders fair comparison.
@@ 6.2 Performance Under Homogeneous Models
To assess the effectiveness of FedMLAC under homogeneous model conditions, we evaluate its performance against three representative FL baselines: FedAvg, FedProx, and FedOPT. In each communication round, only 20% of clients are randomly selected to participate, except for the IEMOCAP dataset, where all clients are activated due to its small client population. Unlike FedMLAC, which maintains personalized local models and exchanges lightweight Plug-in models with the server, baseline methods directly share full client models for aggregation. All methods are evaluated under identical training protocols, and for fair comparison, we report the averaged performance of all local client models at the end of training. This ensures that the comparison reflects the collective quality of personalized models across the federation, rather than relying solely on a global proxy model.
@@
FedAvg FedProx FedOPT FedMLAC
GSC 524 (20%) 80.19$\pm$0.53 80.43$\pm$0.47 83.31$\pm$0.24 84.59$\pm$0.19
IEMOCAP 10 (100%) 48.54$\pm$6.59 48.89$\pm$6.39 50.01$\pm$6.77 51.97$\pm$5.81
CREMA-D 18 (20%) 60.84$\pm$3.89 61.02$\pm$3.71 58.47$\pm$4.34 63.04$\pm$3.42
@@
10 (20%) 54.64$\pm$4.40 54.89$\pm$3.54 53.03$\pm$4.93 56.95$\pm$3.65
As shown in Table 4, FedMLAC consistently outperforms all baseline methods across datasets and varying degrees of data heterogeneity. On Google Speech Commands, a large-scale and relatively balanced dataset, FedMLAC surpasses FedOPT by 1.28%, demonstrating its ability to effectively integrate global knowledge even under limited client participation. For IEMOCAP, which poses challenges due to high inter-speaker variability and data scarcity, FedMLAC yields up to a 3.43% improvement over FedAvg and 3.08% over FedProx, highlighting its robustness in low-resource, high-variance environments.
On the CREMA-D dataset, despite reduced client activation, FedMLAC achieves superior performance, reflecting its resilience to partial participation and speaker-specific skew. In the Urban Sound experiments, where the Dirichlet parameter $\alpha$ controls the degree of heterogeneity of federated data distributions, all methods show improved accuracy as $\alpha$ increases. However, FedMLAC consistently maintains a clear advantage across all heterogeneity levels, with accuracy rising from 52.86% at $\alpha=0.1$ to 56.95% at $\alpha=1.0$. These results collectively underscore the effectiveness of FedMLAC’s mutual learning mechanism and layer-wise robust aggregation in handling both personalization and non-IID challenges in FedAC.
@@ 6.3 Performance Under Heterogeneous Models
In this subsection, we assess the performance of FedMLAC where clients employ heterogeneous local model architectures, a realistic setting in edge computing and cross-device FL. Since conventional FL baselines such as FedAvg, FedProx, and FedOPT require model consistency across clients, they are excluded from this comparison. Instead, we evaluate FedMLAC against its ablated variant (FedMLAC w/o LPA), where the proposed LPA is disabled. All remaining experimental conditions, including client participation ratio and training protocol, are consistent with those in Section 6.2 to ensure fair comparison.
Figure 2: Performance comparison between FedMLAC and FedMLAC without LPA under heterogeneous model settings across four AC datasets. Each bar represents the average accuracy across all clients, with error bars indicating standard deviation. “GSC” denotes Google Speech Commands, and “US” refers to Urban Sound.
As shown in Figure 2, FedMLAC consistently outperforms its variant without LPA across all tested AC datasets and non-IID configurations. On Google Speech Commands, FedMLAC improves accuracy by 1.75%, indicating stable global knowledge transfer even under architectural diversity. In low-resource emotion datasets like IEMOCAP and CREMA-D, where speaker-induced variation is prominent, FedMLAC yields 2.04% and 5.23% accuracy improvements, respectively, suggesting superior robustness against speaker and model shifts. Moreover, on Urban Sound, where Dirichlet-based non-IID data distribution is introduced, FedMLAC achieves steady gains across all levels of heterogeneity. Notably, under the most skewed setting ($\alpha=0.1$), it outperforms its ablated version by 3.82%, demonstrating that LPA plays a key role in filtering noisy updates and stabilizing aggregation under model and data heterogeneity. These results highlight the essential contribution of LPA in achieving generalizable and resilient performance in heterogeneous FedAC systems.
@@ 6.4 Performance Under Varying Active Ratio
FedAvg FedProx FedOPT FedMLAC
GSC 262 (10%) 78.14$\pm$0.52 78.30$\pm$0.49 81.92$\pm$0.22 82.81$\pm$0.17
524 (20%) 80.19$\pm$0.53 80.43$\pm$0.47 83.31$\pm$0.24 84.59$\pm$0.19
1310 (50%) 81.73$\pm$0.58 81.95$\pm$0.35 83.82$\pm$0.31 85.10$\pm$0.21
2618 (100%) 82.58$\pm$0.46 82.73$\pm$0.42 84.84$\pm$0.19 85.76$\pm$0.13
CREMA-D 9 (10%) 59.87$\pm$4.02 60.10$\pm$3.78 58.02$\pm$4.23 61.78$\pm$3.62
18 (20%) 60.84$\pm$3.89 61.02$\pm$3.71 58.47$\pm$4.34 63.04$\pm$3.42
45 (50%) 61.72$\pm$3.75 61.85$\pm$4.16 59.52$\pm$4.05 63.93$\pm$2.24
91 (100%) 62.57$\pm$3.16 62.45$\pm$3.50 60.20$\pm$3.90 64.79$\pm$2.08
Urban Sound 5 (10%) 50.65$\pm$5.21 51.18$\pm$5.23 49.12$\pm$5.89 52.74$\pm$4.26
10 (20%) 53.43$\pm$4.22 53.61$\pm$4.01 51.77$\pm$5.47 55.61$\pm$3.83
25 (50%) 53.52$\pm$5.02 54.29$\pm$3.83 52.35$\pm$5.26 57.15$\pm$5.24
50 (100%) 55.21$\pm$3.96 55.47$\pm$4.74 53.12$\pm$4.41 58.30$\pm$3.59

### v2

@@ 4.4 Discussion
The mutual learning mechanism, via KL divergence in Equations (8) and  (9), aligns local models with global knowledge, reducing the heterogeneity term $\sigma^{2}$. This ensures that local updates contribute effectively to the global objective, even with non-IID data. LPA’s pruning strategy bounds $\delta^{2}$ by excluding anomalous updates, enhancing robustness against data poisoning. By focusing on trusted subsets, LPA mitigates the impact of corrupted or outlier client updates, ensuring stable aggregation.
The $\mathcal{O}(1/\sqrt{T})$ rate is standard for non-convex FL. The error terms reflect challenges inherent to distributed optimization. Limitations include the non-convexity assumption and fixed LPA parameters $v_{h}$, $v_{l}$. Future work could explore adaptive pruning or convex settings to further optimize convergence.
@@ 5 Experiments
In this section, we conduct comprehensive experiments to evaluate the effectiveness of the proposed FedMLAC framework. Our evaluation spans four benchmark datasets covering both speech and non-speech audio classification tasks. We assess FedMLAC in terms of classification accuracy, robustness to noisy and heterogeneous data, and personalization capability, comparing it against four representative state-of-the-art FL baselines.
5.1 Datasets and Models
@@
Datasets. We evaluate FedMLAC on four publicly available datasets widely adopted in federated audio classification (FedAC), as summarized in Table 2. (1) Google Speech Commands (GSC) [37] is a large-scale benchmark for keyword spotting, comprising over 100,000 short utterances covering 35 command words. (2) Urban Sound [38] includes environmental audio samples spanning 10 everyday sound event classes, such as sirens, drilling, and dog barking. (3) IEMOCAP [39] is a multimodal dataset for emotion recognition. In this work, we focus on its audio-only component, selecting four commonly used emotions from the improvised sessions. (4) CREMA-D [40] features speech clips performed by actors under emotional prompts. We consider the audio channel and select four target emotions consistent with prior studies. To ensure fair comparison across datasets, we apply consistent preprocessing strategies including silence trimming, segmentation, and acoustic feature extraction (e.g., Mel spectrograms or pretrained Autoregressive Predictive Coding (APC) embeddings), following the protocol proposed in [41].
Data Partition. To simulate realistic client distributions in federated settings, we adopt dataset-specific non-IID partitioning strategies. For the GSC and CREMA-D datasets, we assign samples to clients based on speaker identity, ensuring that each client contains utterances from a single speaker. This setup reflects the naturally occurring speaker-level data separation commonly found in speech-based federated systems. In the case of IEMOCAP, we divide data by actor identity, with each client corresponding to the recordings of one actor. This preserves speaker consistency within clients while exposing variation across clients. For the Urban Sound dataset, we follow a label distribution skewing strategy using the Dirichlet allocation method, as commonly adopted in FL studies [42]. Specifically, we vary the Dirichlet concentration parameter $\alpha\in\{0.1,0.3,0.5,1.0\}$ to control the degree of data heterogeneity. Smaller values of $\alpha$ lead to greater data heterogeneity across clients, whereas larger values result in more balanced and uniform data distributions.
AC Models. To support AC in both homogeneous and heterogeneous FL settings, we construct a suite of Convolutional Recurrent Neural Network (CRNN) architectures with varying levels of complexity as listed in Table 3. All models share a consistent input-output format and are compatible with the same audio preprocessing pipeline, ensuring fair performance comparison across settings.
@@
5.2 Evaluation and Implementation Details
Baselines. We compare our method with three widely used FL algorithms: FedAvg [35], FedProx [43], and FedOPT [44]. These are strong general-purpose baselines commonly applied across diverse FL tasks, including AC. We also include FedKAD [34], a recent KD-based approach designed to handle both data and model heterogeneity in FedAC. To ensure a fair comparison with our FedMLAC framework, we modify FedKAD by removing its reliance on public datasets. Specifically, we disable the use of global logits during knowledge distillation and restrict communication to class-wise aggregated feature maps only. For other FedAC methods listed in Table 1, fair comparison is not feasible due to the lack of standardized implementations and their dependence on incompatible datasets or architectures.
Metrics. We evaluate performance on the GSC dataset using classification accuracy, averaged over five independent runs with different random seeds. For the other three datasets, we adopt the F1 score as the evaluation metric, reporting the mean and standard deviation across all runs. F1 score is used to account for class imbalance, which is common in emotion recognition and sound event classification tasks.
@@ 5.3 Performance Under Homogeneous Models
To assess the effectiveness of FedMLAC under homogeneous model conditions, we evaluate its performance against four representative FL baselines: FedAvg, FedProx, FedOPT, and FedKAD. In each communication round, only 20% of clients are randomly selected to participate, except for the IEMOCAP dataset, where all clients are activated due to its small client population. As shown in Table 4, FedMLAC consistently outperforms all baseline methods across datasets and varying degrees of data heterogeneity. On GSC, a large-scale and relatively balanced dataset, FedMLAC surpasses FedOPT by 1.28% and improves upon FedKAD by 1.96%, demonstrating its ability to effectively integrate global knowledge even under limited client participation. Notably, while FedKAD shows stronger performance than FedAvg and FedProx, it still lags behind FedMLAC, suggesting that its distillation-based alignment alone is insufficient in capturing global generalization. For IEMOCAP, which poses challenges due to high inter-speaker variability and data scarcity, FedMLAC yields up to a 3.43% improvement over FedAvg and 3.08% over FedProx, and outperforms FedKAD by 1.49%, highlighting its robustness in low-resource, high-variance environments.
@@
FedAvg [35] FedProx [43] FedOPT [44] FedKAD [34] FedMLAC
GSC 524 (20%) 80.19$\pm$0.53 80.43$\pm$0.47 83.31$\pm$0.24 82.63$\pm$0.39 84.59$\pm$0.19
IEMOCAP 10 (100%) 48.54$\pm$6.59 48.89$\pm$6.39 50.01$\pm$6.77 50.48$\pm$5.35 51.97$\pm$5.81
CREMA-D 18 (20%) 60.84$\pm$3.89 61.02$\pm$3.71 58.47$\pm$4.34 62.17$\pm$4.08 63.04$\pm$3.42
@@
10 (20%) 54.64$\pm$4.40 54.89$\pm$3.54 53.03$\pm$4.93 54.95$\pm$5.10 56.95$\pm$3.65
On the CREMA-D dataset, despite reduced client activation, FedMLAC achieves superior performance, reflecting its resilience to partial participation and speaker-specific skew. Compared to FedKAD, which already outperforms FedOPT and FedProx on this task, FedMLAC still obtains a 0.87% gain, demonstrating the advantage of its mutual learning design. In the Urban Sound experiments, where the Dirichlet parameter $\alpha$ controls the degree of heterogeneity of federated data distributions, all methods show improved accuracy as $\alpha$ increases. However, FedMLAC consistently maintains a clear advantage across all heterogeneity levels, with accuracy rising from 52.86% at $\alpha=0.1$ to 56.95% at $\alpha=1.0$. Although FedKAD also shows competitive performance and benefits from increased $\alpha$, it is consistently outperformed by FedMLAC with margins ranging from 0.94% to 2.01%. These results collectively underscore the effectiveness of FedMLAC’s mutual learning mechanism and layer-wise robust aggregation in handling both personalization and non-IID challenges in FedAC.
@@ 5.4 Performance Under Heterogeneous Models
In this subsection, we assess the performance of FedKAD and FedMLAC in scenarios where clients employ heterogeneous local model architectures, which is a realistic setting in edge computing and cross-device FL. Since conventional FL baselines such as FedAvg, FedProx, and FedOPT require model consistency across clients, they are excluded from this comparison. Instead, we evaluate FedMLAC against two alternatives: FedKAD, a recent distillation-based approach designed for model heterogeneity, and an ablated variant of our method (FedMLAC w/o LPA), where the proposed LPA is disabled. All remaining experimental conditions, including client participation ratio and training protocol, are kept consistent with those in Section 5.3 to ensure fair comparison.
Figure 2: Performance comparison among FedKAD, FedMLAC without LPA, and FedMLAC under heterogeneous model settings across four AC datasets. Each bar represents the average accuracy across all clients, with error bars indicating standard deviation. “GSC” denotes Google Speech Commands, and “US” refers to Urban Sound.
As shown in Figure 2, under heterogeneous model settings, FedMLAC consistently outperforms both its ablated variant without LPA and the KD-based baseline FedKAD across all AC datasets and non-IID configurations. Specifically, it achieves the highest accuracy across all cases, improving over its variant by up to 5.23% on CREMA-D and over FedKAD by as much as 4.45%. On GSC, FedMLAC reaches an accuracy of nearly 80%, surpassing other methods with a margin of 1.75%. For Urban Sound across different heterogeneity levels ($\alpha=0.1$, 0.30.30.3, and 0.50.50.5), FedMLAC steadily leads with up to 3.82% gains, demonstrating stable and superior performance. These results highlight FedMLAC’s robustness to model and data heterogeneity, and emphasize the critical role of LPA in filtering noisy updates and enhancing aggregation stability under challenging FedAC conditions.
@@ 5.5 Performance Under Varying Active Ratio
FedAvg [35] FedProx [43] FedOPT [44] FedKAD [34] FedMLAC
GSC 262 (10%) 78.14$\pm$0.52 78.30$\pm$0.49 81.92$\pm$0.22 81.69$\pm$0.33 82.81$\pm$0.17
524 (20%) 80.19$\pm$0.53 80.43$\pm$0.47 83.31$\pm$0.24 82.63$\pm$0.39 84.59$\pm$0.19
1310 (50%) 81.73$\pm$0.58 81.95$\pm$0.35 83.82$\pm$0.31 83.92$\pm$0.29 85.10$\pm$0.21
2618 (100%) 82.58$\pm$0.46 82.73$\pm$0.42 84.84$\pm$0.19 84.71$\pm$0.28 85.76$\pm$0.13
CREMA-D 9 (10%) 59.87$\pm$4.02 60.10$\pm$3.78 58.02$\pm$4.23 61.14$\pm$3.65 61.78$\pm$3.62
18 (20%) 60.84$\pm$3.89 61.02$\pm$3.71 58.47$\pm$4.34 62.17$\pm$4.08 63.04$\pm$3.42
45 (50%) 61.72$\pm$3.75 61.85$\pm$4.16 59.52$\pm$4.05 62.54$\pm$3.79 63.93$\pm$2.24
91 (100%) 62.57$\pm$3.16 62.45$\pm$3.50 60.20$\pm$3.90 63.31$\pm$4.23 64.79$\pm$2.08
Urban Sound 5 (10%) 50.65$\pm$5.21 51.18$\pm$5.23 49.12$\pm$5.89 51.25$\pm$5.04 52.74$\pm$4.26
10 (20%) 53.43$\pm$4.22 53.61$\pm$4.01 51.77$\pm$5.47 53.88$\pm$5.39 55.61$\pm$3.83
25 (50%) 53.52$\pm$5.02 54.29$\pm$3.83 52.35$\pm$5.26 54.65$\pm$4.51 57.15$\pm$5.24
50 (100%) 55.21$\pm$3.96 55.47$\pm$4.74 53.12$\pm$4.41 55.97$\pm$4.87 58.30$\pm$3.59